kiarafranklin13:

During play, a ball with a diameter of 3.00 inches splits exactly in half. What is the volume of each portion of the ball?

psirockin2:

V = 4/3(pi)r^3 , r = 3/2 = 1.5 V = (4/3)(3.14)(1.5^3) 14.10 divide by 2 for 1/2 of the ball. V/2 = 14.10 = 7.05 7.05 in^3

Falconmaster:

113.1 iv think

Falconmaster:

THANKS ALOT GOOGLE IM WRONG BECAUSE OF YOU

kiarafranklin13:

thx